Problem 15-3A Source documents, journal entries, and accounts in job order costing LO P1, P2, P3 [The following information aProblem 15-3A Part 1 Required: 1. Prepare a job cost sheet for each job worked on during the month. Job No. 136 Job No. 137 JProblem 15-3A Part 2 2. Prepare journal entries to record the events and transactions a through i. View transaction list JourProblem 15-3A Part 3 3. Prepare T-accounts for each of the following general ledger accounts, each of which started the monthProblem 15-3A Part 4 4. Prepare a report showing the total cost of each job in process and prove that the sum of their costs
Answer #1

1)

Job No. 136 Job No. 137 Job No. 138 Job No. 139 Job No. 140
Materials $49000 $32500 $19400 $23200 $7200
Labor 12200 10700 37900 39000 3800
Overhead (12200*2)= 24400 (10700*2)= 21400 (37900*2)= 75800 (39000*2)= 78000 (3800*2)= 7600
Total cost $85600 $64600 $133100 $140200 $18600

2)

Transaction General Journal Debit Credit
a Raw materials inventory $240000
Account payable $240000
(To record raw materials purchased on account)
b Work in process inventory $131300
Factory overhead $21000
Raw materials inventory $152300
(To record raw materials used)
c Factory overhead $15750
Cash $15750
(To record cash paid)
d Work in process inventory $103600
Factory overhead $26000
Cash $129600
(To record labor used)
e Work in process inventory ($24400+75800+78000) $178200
Factory overhead $178200
(To record factory overhead applied of Jobs 136, 138 and 139)
f Finished goods inventory ($85600+133100+140200) $358900
Work in process inventory $358900
(To record transferred of Job 136, 138 and 139 to finished goods)
g(1) Account receivable $540000
Sales $540000
(To record sales)
g(2) Cost of goods sold ($85600+133100) $218700
Finished goods inventory $218700
(To record cost of goods sold)
h Factory overhead ($69500+38000+11000+36500) $155000
Accumulated depreciation- Factory building $69500
Accumulated depreciation- Factory equipment $38000
Prepaid insurance $11000
Property tax payable $36500
(To record actual factory overhead)
i Work in process inventory ($21400+7600) $29000
Factory overhead $29000
(To record overhead applied of Job 137 and 140)

3)

Raw materials Inventory Work in process Inventory
Beg. bal. $0 Beg. bal. $0
a 240000 b 131300 358900 f
152300 b d 103600
e 178200
End. bal. $87700 i 29000
End. bal. 83200
Factory overhead Finished goods Inventory
b 21000 Beg. bal. $0
c 15750 178200 e f 358900 218700 g
d 26000 29000 i
h 155000 End. bal. 140200
End. bal. 10550
Cost of Goods Sold
g 218700
End. bal. 218700

4)

Report of Job Costs
Work in process Inventory
Job 137 $64600
Job 140 18600
Balance $83200
Finished goods Inventory
Job 139 $140200
Balance $140200
Cost of Goods Sold
Job 136 $85600
Job 138 133100
Balance $218700